OpenStudy (anonymous):

Solve for x and y 5/4x-3/4y=9/4 5/3y-7/3x=15/3

OpenStudy (anonymous):

use gauss-jordan elimination

OpenStudy (anonymous):

i got to 5x-3y=9 -7x+5y=15

OpenStudy (anonymous):

then i multiplied the first one by 5 then the second by 3 to get 25x-15y=45 -21x+15y=45

OpenStudy (anonymous):

that works

OpenStudy (anonymous):

add both equations together to get a solution for x 4x = 90 x = 45 then you can find y.

OpenStudy (anonymous):

when i divided 90 by 4 i got 22.5...